### v1

@@ 4.5 Incorporating Bounding Box Information
To assess the impact of spatial information, we incorporated bounding box coordinates into the prompts, denoted as Reph-Reph-bbox in Table 4. Each Textract-extracted text element in the prompt was annotated with bounding box coordinates, enabling the model to reference spatial context. In this configuration, the model was specifically fine-tuned to produce more complex JSON outputs that include not only the answer but also a comprehensive list of all locations where the extracted value appears in the document. While this approach provided richer spatial awareness, the requirement to generate more structured outputs introduced additional complexity that led to increased parsing errors.
To address these parsing challenges, we implemented the Reph-Reph-bbox w/regex configuration, which introduced a regex-based post-processing step. When the model’s structured JSON output was not parsable due to format inconsistencies or generation errors, the regex extraction mechanism served as a fallback solution to retrieve the target value, effectively maintaining the benefits of spatial information while mitigating the impact of parsing failures.
@@ 4.6 Research Question answers
  • RQ1 - Dataset Unification: By standardizing data from various sources (e.g., receipts, invoices, forms) into a consistent Question-Answering format, models are exposed to a wide range of document layouts and content types, enhancing their training efficiency. This unification significantly streamlines the fine-tuning process, making it easier to handle diverse document sources. Moreover, fine-tuned models show significant improvements over instruct models, quantitatively confirming that exposure to varied document formats and layouts enhances the model’s ability to extract information.
  • RQ2 - Question Formulation Impact: The study revealed significant insights into how different question formulation strategies affect document comprehension and information extraction. The Template-Template configuration demonstrated superior performance by leveraging structured, consistent question patterns. However, the Rephrased-Rephrased configuration emerged as a particularly robust solution, maintaining high ANLS* scores (e.g., 99.8 on FATURA, 96.4 on VRDU-Ad) while achieving 0% parsing errors across most datasets. This configuration showed remarkable versatility in handling both template-based and natural language queries. Notably, the Template-Rephrased setup performed least effectively, highlighting the challenges in transitioning from template-trained models to complex question structures.
  • RQ3 - Layout Information: The incorporation of spatial information in prompts yielded measurable improvements in model performance. The Reph-Reph-bbox configuration achieved the highest weighted average ANLS* (91.6) across all datasets, demonstrating consistent improvements over configurations without spatial information. Notable gains were observed in complex document understanding tasks, with ANLS* scores increasing to 71.2 on XFUND and 83.0 on SP-VQA. While the initial implementation showed increased parsing errors, the addition of regex-based post-processing (Reph-Reph-bbox w/regex) successfully maintained high performance while reducing error rates to competitive levels (1.53% weighted average).
@@ 5 Conclusions
The paper addresses the growing need for evaluating LLMs in Document AI tasks by proposing a unified dataset designed for document Question Answering taking into account the position of answers’ text in the document. Baseline experiments using open-weight LLMs demonstrate the challenges of applying generic models to specialized Document AI tasks; the performance of instruct models reveal clear limitations in generating correct and well-structured answers.
@@
In future work we aim to explore various methods for incorporating bounding boxes into prompts to better capture the spatial structure of documents and evaluate open-weight multimodal LLMs specifically designed to handle both textual and visual information. The constructed dataset and the experimental results provide a solid foundation for future research in Document QA. Fine-tuning models with enriched prompts has shown promising improvements.
@@ Appendix A Dataset statistics
Figure 6 and Figure 7 provide an overview of the dataset construction, showing how many documents and related questions from the various source datasets contribute to the overall dataset.
@@
In Figure 8 and Table 5 you can observe the distribution of the languages in which the questions in the dataset are posed. The only questions, along with their respective documents, that are not in English are those formulated on XFUND, and thus they represent a clear minority compared to the total count.
@@
Question Type Count Percentage (%)
@@
Question Type Count Percentage (%)
@@
Total amount after tax and discount 7992 7.80
@@
Total amount before tax and discount 6753 6.59
@@
Question Type Count Percentage (%)
@@
Spending Annually in British Pounds 66 0.74
@@
Income Annually in British Pounds 61 0.69
@@
Question Type Count Percentage (%)
@@
Question Type Count Percentage (%)
@@
Figure 8: Language distribution of questions
For Deepform, there are only five fields for which questions have been constructed, as can be observed in Table 6 and Figure 9. The main fields present are the total cost incurred for the advertisement and the name of the advertiser. The fields Flight From and Flight To are date values that represent the start and end days of the spot’s transmission.
Figure 9: Deepform question distribution
Regarding FATURA, the range of extracted fields is much broader compared to the previous Deepform, as visible in Figure 10 and Table 7. With 50 different layouts within FATURA, not all documents contain the same fields, which is the reason for the significant differences in frequencies among some fields. It is notable that the field with the most questions is Date of purchase (9800), while the least frequent is Total amount to be paid (685).
Figure 10: FATURA question distribution
Regarding the Kleister Charity dataset, the range of extracted fields is relatively narrow compared to the FATURA dataset. As shown in Table 8 and Figure 11, the dataset primarily focuses on extracting information such as the Charity Name, Charity Number, Address Post Town, Address Postcode, and Address Street Line. The fields with the fewest questions are Spending Annually in British Pounds and Income Annually in British Pounds, indicating that financial details are less frequently extracted from this dataset.
Figure 11: Kleister Charity question distribution
The Kleister NDA dataset, as detailed in Table and Figure 12, contains questions across a very limited set of fields: Jurisdiction, Party, Term, and Effective Date. The field with the most questions is Jurisdiction, followed by Party, while the Effective Date field has only a single question.
Figure 12: Kleister NDA question distribution
The VRDU Ad Buy Form dataset, as shown in Table 10 and Figure 13, contains a broader range of fields compared to the previous datasets. The fields with the most questions are Program Start Date, Channel, and Program End Date. In contrast, the field with the fewest questions is Program Description, with only 6 questions.
Figure 13: VRDU Ad Buy Form question distribution
The VRDU Registration Form dataset, as detailed in Table 11 and Figure 14, contains questions across 6 different fields. The fields with the most questions are Registration Number and Registrant Name, while the field with the fewest questions is Signer Title.
Question Type Count Percentage(%)
@@
Figure 14: VRDU Registration Form question distribution
@@
Split Documents Questions % Documents % Questions
@@
Appendix B Dataset examples
@@
What is Advertiser? Who is the advertiser? Jordan, Jonathan 15
What is Gross Amount? What is the value of the Gross Amount? \$10,500.00 15
What is Address Post Town? What is the post town of the address? Stoke-on-Trent 16
What is Address Postcode? What is the postal code of the address? ST4 8AW 16
What is Address Street Line? What is the value of the Address Street Line? 28 Greenway 16
What is Charity Name? What is the name of the charity? Lucas’ Legacy - Childhood Brain Tumour Research 16
What is Charity Number? What is the charity number? 1167650 16
What is Jurisdiction? In which state is the company registered? Delaware 17
What is Party? What is the name of the company? Cisco Systems, Inc., 17
What is Contract ID? What is the contract ID number? 711207 18
What is the Product? What is the name of the product being advertised? Q42020 Broadcast 18
What is Property? What is the property name? KXLF 18
What is Agency? Who is the advertising agency? Left Hook Communications 18
@@
18
What is Gross Amount? What is the value of the gross amount? \$3,020.00 18
What is Sub Amount for M-F 530-7am News M-F 530-7am News? What is the value for the ’Sub Amount’ key for ’M-F 530-7am News’? \$100.00 18
What is the Channel for M-F 530-7am News M-F 530-7am News? What is the value of the ’Channel’ for the ’530-7am News’ broadcasted from Monday to Friday? All 18
What is Program Start Date for M-F 530-7am News M-F 530-7am News? What is the start date for the M-F 530-7am News program? 10/06/20 18
What is the Program End Date for M-F 530-7am News M-F 530-7am News? What is the end date for the program ’M-F 530-7am News’? 10/12/20 18
What is Registrant Name? What is the name of the registrant? KOREA TRADE PROMOTION CENTER 19
What is Registration Number? What is the registration number for the company? 1619 19
What is Signer Title? What is the signer’s title? DEPUTY DIRECTOR 19
First bubble in the HPA Axis? First bubble in the HPA Axis? Hypothalamus 20
What does CORT stand for in this document? What does CORT stand for in this document? Cortisol 20
Where does cortisol go after it is sent from the adrenal cortex? Where does cortisol go after it is sent from the adrenal cortex? Hypothalamus 20
@@
21
What is Date of purchase? When was the purchase date? Invoice Date: 30-Oct-1998 21
What is Due date? What is the due date? Due Date : 24-May-2020 21
What is Purchase order number? What is the purchase order number value? PO Number :72 21
What is Seller Address? What is the seller’s address? Address:05866 Velazquez Mount North Diane, NJ 20651 US 21
What is Total amount before tax and discount? What is the value of the total amount before tax and discount? SUB_TOTAL : 293.47 \$ 21
What is Tax? What is the tax amount? TAX:VAT (5.69%): 16.70 \$ 21
What is Title? What is the key for the title information? TAX INVOICE 21
What is Total amount to be paid? What is the value of the total amount to be paid? BALANCE_DUE : 305.39 \$ 21
What is MANUFACTURER:? What is the value of the manufacturer? R. J. REYNOLDS 22
What is BRAND NAME:? What is the value of the brand name? CARDINAL CIGARETTES (11 PACKINGS) 22
What is OTHER INFORMATION:? What is the value of OTHER INFORMATION? SEE ATTACHMENT 22
to whom is this letter written to? to whom is this letter written to? Mr. Rionda 23
when is the letter dated ? when is the letter dated ? October 18, 1940, 23
what is the auth. no. mentioned in the given form ? what is the auth. no. mentioned in the given form ? 5754 24
what is the value of percent per account as mentioned in the given form ? what is the value of percent per account as mentioned in the given form ? 50.06 24
what is the emp. no. mentioned in the given form ? what is the emp. no. mentioned in the given form ? 483378 24
what is the employee name mentioned in the given form ? what is the employee name mentioned in the given form ? IRENE KARL 24
what is the value of amount authorized per account ? what is the value of amount authorized per account ? 292.00 24
Qual è Cognome? Qual è Cognome? ANNI 25
Qual è Nome? Qual è Nome? GIACCOMO 25
Qual è Data Nascita? Qual è Data Nascita? 12/02/1988 25
Qual è Data? Qual è Data? 19/12/2020 25
Qual è Ora? Qual è Ora? 14:00 25

### v2

@@ 4.5 Incorporating bounding box information
To assess the impact of spatial information, we incorporated bounding box coordinates into the prompts, denoted as Reph.-Reph.-bbox in Table 6. Each Textract-extracted text element in the prompt was annotated with bounding box coordinates, enabling the model to reference spatial context. In this configuration, the model was specifically fine-tuned to produce more complex JSON outputs that include not only the answer, but also a comprehensive list of all locations where the extracted value appears in the document. While this approach provided richer spatial awareness, the requirement to generate more structured outputs introduced additional complexity that led to increased parsing errors.
To address these parsing challenges, we implemented the Reph.-Reph.-bbox w/regex configuration, which introduced a regex-based post-processing step. When the model’s structured JSON output was not parsable due to format inconsistencies or generation errors, the regex extraction mechanism served as a fallback solution to retrieve the target value, effectively maintaining the benefits of spatial information while mitigating the impact of parsing failures.
4.6 Fine-Tunings
Table 7 presents the fine-tuning results for a subset of the models listed in Table 5. In these experiments, models were trained and tested using rephrased questions. For Mistral, the prompt included both the page content and bounding boxes (corresponding to the Reph.-Reph.-bbox row in the Table 6), whereas the two Qwen models received only the page image and the rephrased question as input. It is important to noice that these results are not directly comparable across models, as the fine-tuning settings for Mistral and Qwen differ significantly (details on fine-tuning settings can be found in the Supplementary Material.) Once again, these results are not intended for direct comparison, but rather to provide a baseline for the expected performance on this dataset after a fine-tuning of various models.
Model
Deepform
DUDE
FATURA
FUNSD
XFUND
SP-VQA
Kl. Char.
Kl. NDA
MP-VQA
VRDU Ad
VRDU Reg
W. Avg.
Mistral-7B-v0.3 97.7 72.1 99.3 74.7 70.3 83.0 92.9 61.9 75.8 96.1 96.8 91.3
0.80% 0.38% 4.50% 4.29% 0.98% 0.00% 0.03% 0.00% 0.35% 0.26% 0.00% 1.53%
Qwen2-VL-2B 88.0 58.9 97.2 72.9 62.4 79.7 74.1 45.6 66.1 75.9 79.9 82.1
0.05% 0.00% 0.30% 0.61% 0.92% 0.00% 0.43% 0.49% 0.20% 0.78% 0.00% 0.26%
Qwen2-VL-7B 90.6 71.0 98.5 75.6 68.5 84.2 82.8 40.9 74.9 81.2 86.0 86.6
0.01% 0.76% 0.00% 0.61% 0.43% 0.00% 0.43% 0.00% 0.28% 0.00% 0.51% 0.13%
Table 7: Fine-tunings. ANLS* scores and JSON parsing error percentages across datasets for each fine-tuned model on our test split. ANLS* scores measure accuracy in answering document questions, while the bottom value in each cell shows JSON parsing errors, indicating output consistency. The ”W. Avg” column provides a weighted average across datasets, with bold and underlined values marking the top two scores per dataset.
4.7 Answer localization
In this article, we investigate the unique feature of BoundingDocs, bounding boxes, specifically to evaluate whether their inclusion can enhance document understanding. This can be achieved either by incorporating bounding box information as an additional input in textual prompts for off-the-shelf models, or by leveraging it for fine-tuning. Additionally, BoundingDocs can serve as a benchmark for assessing a model’s ability to accurately predict the location of an answer within a page. Our experiments, reported in [chen2025reliableinterpretabledocumentquestion], show that Claude 4 Sonnet achieves an average IoU of 0.031 on the BoundingDocs test set, while Qwen2.5-VL-7B reaches 0.048. These results demonstrate that, although off-the-shelf models perform reasonably well in retrieving information in a zero-shot setting, their capacity to indicate the precise location of the answer is effectively nonexistent, which significantly undermines the interpretability of their responses.
@@ 4.8 Research question answers
  • RQ1 - Dataset Unification: By standardizing data from various sources (e.g., receipts, invoices, forms) into a consistent Question-Answering format, models are exposed to a wide range of document layouts and content types, enhancing their training efficiency. This unification significantly streamlines the fine-tuning process, making it easier to handle diverse document sources. Moreover, fine-tuned models show significant improvements over instruct models, quantitatively confirming that exposure to varied document formats and layouts enhances the model’s ability to extract information, as can be seen comparing the overall results in Table 5 and Table 7.
  • RQ2 - Question Formulation Impact: The study revealed significant insights into how different question formulation strategies affect document comprehension and information extraction. As reported in Table 6, the Template-Template configuration demonstrated superior performance by leveraging structured, consistent question patterns. However, the Rephrased-Rephrased configuration emerged as a particularly robust solution, maintaining high ANLS* scores (e.g., 99.8 on FATURA, 96.4 on VRDU-Ad) while achieving 0% parsing errors across most datasets. Crucially, unlike other configurations, its performance remained stable regardless of whether the input followed a template or an augmented version, indicating that the model had developed a greater adaptability to diverse question structures. Notably, the Template-Rephrased setup performed least effectively, highlighting the challenges in transitioning from template-trained models to complex question structures.
  • RQ3 - Layout Information: The incorporation of spatial information in prompts yielded measurable improvements in model performance, as reported in Table 7. The Reph-Reph-bbox configuration achieved the highest weighted average ANLS* (91.6) across all datasets, demonstrating consistent improvements over configurations without spatial information. Notable gains were observed in complex document understanding tasks, with ANLS* scores increasing to 71.2 on XFUND and 83.0 on SP-VQA. While the initial implementation showed increased parsing errors, the addition of regex-based post-processing (Reph-Reph-bbox w/regex) successfully maintained high performance while reducing error rates to competitive levels (1.53% weighted average).
@@ 5 Conclusions
The paper addresses the growing need for evaluating LLMs in Document AI tasks by proposing a unified dataset designed for document QA taking into account the position of answers’ text in the document. Baseline experiments using off-the-shelf LLMs demonstrate the challenges of applying generic models to specialized Document AI tasks; the performance of instruct models reveal clear limitations in generating correct and well-structured answers.
@@
In future work, we aim to explore a wider range of prompting techniques tailored to different model families. Among the many types of information available in BoundingDocs, a key challenge is determining the optimal combination for effective information extraction. Open questions include whether images can fully replace textual content in prompts, whether bounding boxes remain essential even when using images, and how different modalities interact to enhance model performance. The constructed dataset and the experimental results provide a solid foundation for future research in Document QA. Fine-tuning models with enriched prompts has shown promising improvements.
@@ Appendix A Dataset statistics
Figure 6 provide an overview of the dataset construction, showing how many documents and related questions from the various source datasets contribute to the overall dataset.
@@
In Table 8 you can observe the distribution of the languages in which the questions in the dataset are posed. The only questions, along with their respective documents, that are not in English are those formulated on XFUND, and thus they represent a clear minority compared to the total count.
@@
For Deepform, there are only five fields for which questions have been constructed, as can be observed in Table 9. The main fields present are the total cost incurred for the advertisement and the name of the advertiser. The fields Flight From and Flight To are date values that represent the start and end days of the spot’s transmission.
Question Type Count (%)
@@
Regarding FATURA, the range of extracted fields is much broader compared to the previous Deepform, as visible in Table 10. With 50 different layouts within FATURA, not all documents contain the same fields, which is the reason for the significant differences in frequencies among some fields. It is notable that the field with the most questions is Date of purchase (9800), while the least frequent is Total amount to be paid (685).
Question Type Count (%)
@@
Total amount after tax 7992 7.80
@@
Total amount before tax 6753 6.59
@@
Question Type Count (%)
@@
Spending Annually in £ 66 0.74
@@
Income Annually in £ 61 0.69
@@
Regarding the Kleister Charity dataset, the range of extracted fields is relatively narrow compared to the FATURA dataset. As shown in Table 11, the dataset primarily focuses on extracting information such as the Charity Name, Charity Number, Address Post Town, Address Postcode, and Address Street Line. The fields with the fewest questions are Spending Annually in British Pounds and Income Annually in British Pounds, indicating that financial details are less frequently extracted from this dataset.
The Kleister NDA dataset, as detailed in Table 12, contains questions across a very limited set of fields: Jurisdiction, Party, Term, and Effective Date. The field with the most questions is Jurisdiction, followed by Party, while the Effective Date field has only a single question.
Question Type Count (%)
@@
The VRDU Ad Buy Form dataset, as shown in Table 13, contains a broader range of fields compared to the previous datasets. The fields with the most questions are Program Start Date, Channel, and Program End Date. In contrast, the field with the fewest questions is Program Description, with only 6 questions.
Question Type Count (%)
@@
The VRDU Registration Form dataset, as detailed in Table 14, contains questions across 6 different fields. The fields with the most questions are Registration Number and Registrant Name, while the field with the fewest questions is Signer Title.
Question Type Count (%)
@@
Split Docs Questions % Docs % Questions
@@
Appendix B Questions rephrasing
As outlined in the main text, a key method for generating BoundingDocs is question rewriting using LLMs, specifically Mistral 7B. The model was employed through a few-shot approach, as illustrated in Figure 7, where specific examples of the desired question reformulations were provided. The prompt did not include the answer to the question, as we observed that incorporating the answer in the few-shot examples influenced the question formulation. This led to the generation of questions biased toward facilitating information retrieval, thereby compromising the neutrality of the rewriting process.
Questions rephrasing User: I have a set of questions written in natural language using a simple template. These questions may contain grammatical errors, unnatural phrasing, or lack richness in expression. Your task is to refine them while keeping their original intent. — TASK — - Correct any grammatical mistakes - Improve fluency and clarity - Enrich phrasing to make the question more natural and well-formed — EXAMPLES — - Original: [What is Charity Name?] Rephrased: [What is the name of the Charity?] - Original: [What is the Product?] Rephrased: [What is the name of the product?] — QUESTION TO REPHRASE — - Original: [What is Jurisdiction?] LLM: In which state is the company registered?
Figure 7: Prompt example and LLM’s answer for questions rephrasing.
Appendix C Finetuning details
For our work, we fine-tuned three different models: Mistral-7B-v0.3, Qwen2-VL-2B, and Qwen2-VL-7B. All three models were fine-tuned on the 10% of the train split using QLoRA, but with two different configurations: one for Mistral, which processes only text input, and another for the Qwen2-VL models, which also take document images as input. The fine-tuning process was performed in Python using the HuggingFace SFT trainer. Quantization was achieved using the BitsAndBytes library, and LoRA was applied through the peft, both of which are also Python libraries.
Parameter Mistral-v0.3 Qwen2-VL
load_in_4bit True True
use_double_quant True True
quant_type "nf4" "nf4"
compute_dtype bfloat16 bfloat16
lora_alpha 128 16
lora_dropout 0.05 0.05
r 256 8
bias "none" "none"
target_modules "all-linear" "q_proj"
task_type "CAUSAL_LM" "CAUSAL_LM"
Table 16: Fine-tuning configurations for the Mistral and Qwen2-VL models.
Appendix D Dataset examples
@@
What is Advertiser? Who is the advertiser? Jordan, Jonathan 8
What is Gross Amount? What is the value of the Gross Amount? \$10,500.00 8
What is Address Post Town? What is the post town of the address? Stoke-on-Trent 9
What is Address Postcode? What is the postal code of the address? ST4 8AW 9
What is Address Street Line? What is the value of the Address Street Line? 28 Greenway 9
What is Charity Name? What is the name of the charity? Lucas’ Legacy - Childhood Brain Tumour Research 9
What is Charity Number? What is the charity number? 1167650 9
What is Jurisdiction? In which state is the company registered? Delaware 10
What is Party? What is the name of the company? Cisco Systems, Inc., 10
What is Contract ID? What is the contract ID number? 711207 11
What is the Product? What is the name of the product being advertised? Q42020 Broadcast 11
What is Property? What is the property name? KXLF 11
What is Agency? Who is the advertising agency? Left Hook Communications 11
@@
11
What is Gross Amount? What is the value of the gross amount? \$3,020.00 11
What is Sub Amount for M-F 530-7am News M-F 530-7am News? What is the value for the ’Sub Amount’ key for ’M-F 530-7am News’? \$100.00 11
What is the Channel for M-F 530-7am News M-F 530-7am News? What is the value of the ’Channel’ for the ’530-7am News’ broadcasted from Monday to Friday? All 11
What is Program Start Date for M-F 530-7am News M-F 530-7am News? What is the start date for the M-F 530-7am News program? 10/06/20 11
What is the Program End Date for M-F 530-7am News M-F 530-7am News? What is the end date for the program ’M-F 530-7am News’? 10/12/20 11
What is Registrant Name? What is the name of the registrant? KOREA TRADE PROMOTION CENTER 12
What is Registration Number? What is the registration number for the company? 1619 12
What is Signer Title? What is the signer’s title? DEPUTY DIRECTOR 12
First bubble in the HPA Axis? First bubble in the HPA Axis? Hypothalamus 13
What does CORT stand for in this document? What does CORT stand for in this document? Cortisol 13
Where does cortisol go after it is sent from the adrenal cortex? Where does cortisol go after it is sent from the adrenal cortex? Hypothalamus 13
@@
14
What is Date of purchase? When was the purchase date? Invoice Date: 30-Oct-1998 14
What is Due date? What is the due date? Due Date : 24-May-2020 14
What is Purchase order number? What is the purchase order number value? PO Number :72 14
What is Seller Address? What is the seller’s address? Address:05866 Velazquez Mount North Diane, NJ 20651 US 14
What is Total amount before tax and discount? What is the value of the total amount before tax and discount? SUB_TOTAL : 293.47 \$ 14
What is Tax? What is the tax amount? TAX:VAT (5.69%): 16.70 \$ 14
What is Title? What is the key for the title information? TAX INVOICE 14
What is Total amount to be paid? What is the value of the total amount to be paid? BALANCE_DUE : 305.39 \$ 14
What is MANUFACTURER:? What is the value of the manufacturer? R. J. REYNOLDS 15
What is BRAND NAME:? What is the value of the brand name? CARDINAL CIGARETTES (11 PACKINGS) 15
What is OTHER INFORMATION:? What is the value of OTHER INFORMATION? SEE ATTACHMENT 15
to whom is this letter written to? to whom is this letter written to? Mr. Rionda 16
when is the letter dated ? when is the letter dated ? October 18, 1940, 16
what is the auth. no. mentioned in the given form ? what is the auth. no. mentioned in the given form ? 5754 17
what is the value of percent per account as mentioned in the given form ? what is the value of percent per account as mentioned in the given form ? 50.06 17
what is the emp. no. mentioned in the given form ? what is the emp. no. mentioned in the given form ? 483378 17
what is the employee name mentioned in the given form ? what is the employee name mentioned in the given form ? IRENE KARL 17
what is the value of amount authorized per account ? what is the value of amount authorized per account ? 292.00 17
Qual è Cognome? Qual è Cognome? ANNI 18
Qual è Nome? Qual è Nome? GIACCOMO 18
Qual è Data Nascita? Qual è Data Nascita? 12/02/1988 18
Qual è Data? Qual è Data? 19/12/2020 18
Qual è Ora? Qual è Ora? 14:00 18